# MagicFace: High-Fidelity Facial Expression Editing with Action-Unit Control

### v2

@@ V Impact Statement
VI Limitations and Conclusion
Limitations. Our model may struggle to generate completely the same image when the AU variations are inputted as all zeros, partially because of the inaccuracy of the estimated AU intensity used in training stage. Besides, our model demonstrates lower operational efficiency compared to non-diffusion-based methods due to the use of DDPM.
Conclusion. In this paper, we present MagicFace, a framework capable of editing any portraits with different facial expressions and intensity levels by embedding action units as condition into a Stable-Diffusion. We introduce ID encoder that effectively retains detailed face attributes and achieve efficient facial expression controllability. Our approach provides an user-friendly way to enable fine-grained control over facial expressions with high quality, which outperforms existing face expression editing methods.

### v3

@@ V Impact Statement
VI Discussion and Conclusion
Limitations. Our model may struggle to generate the same image completely when the AU variations are input as all zeros, partially because of the inaccuracy of the estimated AU intensity used in the training stage. Besides, our model demonstrates lower operational efficiency compared to non-diffusion-based methods due to the use of DDPM.
Conclusion. In this paper, we present MagicFace, a framework capable of editing any portraits with different facial expressions and intensity levels by embedding action units as conditions into a Stable-Diffusion. We introduce an ID encoder that effectively retains detailed face attributes and achieves efficient facial expression controllability. Our approach provides a user-friendly way to enable fine-grained control over facial expressions with high quality, which outperforms existing face expression editing methods.
Future Work. While the current work focuses on static AU-based facial expression editing, several promising extensions remain for future research. First, we aim to develop AU-based facial animation by enabling temporally consistent expression control across video frames. This is particularly important because in dynamic scenarios the head pose changes continuously, and the visible background consequently shifts from frame to frame. Addressing this requires incorporating a component that can infer and update the background as head pose and facial geometry evolve over time, thereby avoiding cross-frame inconsistencies and enabling smooth AU-driven facial motion. Second, we intend to scale our approach to high-resolution facial expression editing (e.g., 1024×1024 or higher). Achieving this level of fidelity will require more advanced generative backbones, such as SDXL  [^5] https://stablediffusionxl.com/ or other state-of-the-art diffusion architectures, which can better preserve identity and fine-grained facial details under AU manipulation. These two directions represent promising next steps toward more expressive and versatile facial expression editing.